Write the slope intercept form of the equation of the line that is perpendicula to 3x+9y=12 and contains the point -4,5

Answer by mananth(16946)

3 x + 9 y = 12
Find the slope of this line

9 y = -3 x + 12
Divide by 9
y = - 1/ 3 x + 1 1/3
Compare this equation with y=mx+b, m= slope & b= y intercept
slope m = - 1/3

The slope of a line perpendicular to the above line will be the negative reciprocal 3
Because m1*m2 =-1
The slope of the required line will be 3

m= 3 ,point ( -4 , 5 )
Find b by plugging the values of m & the point in
y=mx+b
5 = -12 + b
b= 17
m= 3
The required equation is y = 3 x + 17
m.ananth@hotmail.ca